# Partial-Label Learning with a Reject Option

### v3

@@ 4.4 Runtime Complexity
Algorithm 1 yields a worst-case runtime complexity of $\mathcal{O}\!\left(dkn+l\left(\min\!\left(2^{k},2^{l}\right)+\max\!\left(k,l\right)\right)\right)$. Since $k$ and $l$ are constant, the nearest-neighbor search dominates. The average total runtime of the search is $\mathcal{O}\!\left(dk\log n\right)$. An in-depth discussion can be found in Appendix A.4.
@@ 5 Experiments
Section 5.1 lists the tested methods, Section 5.2 shows our setup, and Section 5.3 collects our main findings. Additional results and a description of all hyperparameters can be found in Appendix D.
@@ 5.1 Algorithms for Comparison
While many PLL algorithms exist (see Section 3), we focus on methods commonly used in the literature. In total, we consider ten methods: Pl-Knn (Hüllermeier & Beringer, 2005), Pl-Svm (Nguyen & Caruana, 2008), Ipal (Zhang & Yu, 2015), Pl-Ecoc (Zhang et al., 2017), Proden (Lv et al., 2020), Rc (Feng et al., 2020), Cc (Feng et al., 2020), Valen (Xu et al., 2021), Pop (Xu et al., 2023), and Dst-Pll (our proposed method). We choose their parameters as recommended and use the same base models for all of the neural network approaches. As base models, we pick the LeNet architecture (LeCun et al., 1998) for the MNIST datasets and a $d$-300-300-300-$l$ MLP (Werbos, 1974) for all other datasets. For Pl-Knn and our method, we use variational auto-encoders (Kingma & Welling, 2014) to reduce the feature space dimensionality of the MNIST datasets and compute the nearest neighbors on the hidden representations. As our competitors do not provide a reject option, we use a threshold on their model’s confidences, that is, the maximum probability outputs, to evaluate the trade-off between the fraction and accuracy of non-rejected predictions. Appendix D.2 discusses the choices of all hyperparameters in more detail.
@@
Pl-Knn (2005) 81.2 ($\pm$ 13.9) 75.7 ($\pm$ 11.4) 92.2 ($\pm$ 04.1) 84.9 ($\pm$ 07.2) 53.4 ($\pm$ 11.0)
Pl-Svm (2008) 62.3 ($\pm$ 16.3) 42.5 ($\pm$ 17.1) 67.5 ($\pm$ 09.8) 46.8 ($\pm$ 06.9) 39.1 ($\pm$ 09.6)
Ipal (2015) 79.4 ($\pm$ 16.9) 75.6 ($\pm$ 18.1) 92.9 ($\pm$ 04.3) 88.5 ($\pm$ 06.6) 58.7 ($\pm$ 09.8)
Pl-Ecoc (2017) 63.7 ($\pm$ 13.1) 66.5 ($\pm$ 12.1) 64.3 ($\pm$ 14.3) 51.6 ($\pm$ 11.2) 46.2 ($\pm$ 10.2)
Proden (2020) 81.4 ($\pm$ 14.0) 78.1 ($\pm$ 13.2) 93.9 ($\pm$ 04.5) 88.3 ($\pm$ 06.1) 64.0 ($\pm$ 07.9)
Rc (2020) 82.0 ($\pm$ 13.9) 78.3 ($\pm$ 13.6) 94.0 ($\pm$ 04.1) 88.0 ($\pm$ 06.2) 61.5 ($\pm$ 10.0)
Cc (2020) 81.8 ($\pm$ 14.2) 78.5 ($\pm$ 13.8) 94.0 ($\pm$ 04.5) 89.6 ($\pm$ 05.7) 49.0 ($\pm$ 29.6)
Valen (2021) 79.6 ($\pm$ 15.3) 75.3 ($\pm$ 12.8) 91.9 ($\pm$ 04.1) 82.9 ($\pm$ 08.9) 63.3 ($\pm$ 09.5)
Pop (2023) 81.7 ($\pm$ 13.9) 78.0 ($\pm$ 13.1) 93.8 ($\pm$ 04.5) 88.0 ($\pm$ 06.1) 64.0 ($\pm$ 08.2)
Dst-Pll (ours) 80.7 ($\pm$ 14.1) 75.2 ($\pm$ 10.9) 92.0 ($\pm$ 04.2) 84.6 ($\pm$ 07.1) 52.1 ($\pm$ 11.6)
@@
5.2 Experimental Setup
Data. Following the default protocol (Cour et al., 2011; Zhang & Yu, 2015; Lv et al., 2020; Xu et al., 2023), we conduct several experiments using datasets for supervised learning with added artificial noise as well as experiments on real-world partially-labeled data. We repeat all experiments five times to report averages and standard deviations. For the supervised datasets, we use the ecoli (Horton & Nakai, 1996), multiple-features (Duin, 2002), pen-digits (Alpaydin & Alimoglu, 1998), semeion (Buscema & Terzi, 2008), solar-flare (Dodson & Hedeman, 1989), statlog-landsat (Srinivasan, 1993), and theorem datasets (Bridge et al., 2013) from the UCI repository (Bache & Lichman, 2013). Also, we use the MNIST (LeCun et al., 1999), KMNIST (Clanuwat et al., 2018), and FMNIST datasets (Xiao et al., 2017). For the partially labeled data, we use the bird-song (Briggs et al., 2012), flickr (Huiskes & Lew, 2008), yahoo-news (Guillaumin et al., 2010), and msrc-v2 datasets (Liu & Dietterich, 2012).
[…]
5.3 Results
@@
Figure 1 shows the reject trade-off for varying confidence (0 to 1) and $\Delta_{\operatorname{\tilde{m}}}$ (-1 to 1) thresholds on the ecoli and KMNIST datasets with instance-dependent noise as well as on the msrc-v2 real-world dataset. The x-axes show the fractions of predictions that are rejected. The y-axes show the accuracies of predictions that are not rejected. Most methods have monotonical growth: The more predictions are rejected, the more accurate are non-rejected predictions. Also, it is desirable to be close to the top-left corner of the plots as one wants to achieve high accuracy by rejecting as few predictions as possible. Our method provides the significantly best trade-offs compared to competitors (Table 2). These results empirically demonstrate our method’s superiority in reliably rejecting unsure predictions.
@@ 6 Conclusions
We addressed the PLL problem. Current methods offer good predictive performance but even the best algorithms give incorrect predictions. Therefore, we discussed the option to abstain from predictions, that is, to reject individual predictions if unsure. The proposed approach offers the best trade-off between the number and accuracy of non-rejected predictions compared to competitors. Experiments demonstrate the effectiveness of our algorithm. Further, we prove our method’s risk consistency.
Acknowledgments and Disclosure of Funding
This work was supported by the German Research Foundation (DFG) Research Training Group GRK 2153: Energy Status Data - Informatics Methods for its Collection, Analysis and Exploitation.
@@ Appendix A Proofs
This section collects our proofs. The proof of Lemma 4.2 is in Appendix A.1. We prove our main statement, Theorem 4.3, in Appendix A.2. Appendix A.3 contains the proof of Lemma 4.4.
A.1 Proof of Lemma 4.2
@@
A.3 Proof of Lemma 4.4
[…]
A.4 Runtime Analysis of Algorithm 1
Looking at Algorithm 1, we decompose the overall runtime of our approach into $k$-times querying one nearest neighbor and creating its bpa (Lines 2–9), the cost of Yager’s rule (Line 10), as well as extracting predictions in Lines 11-12.
Using the ball-tree data structure (Omohundro, 1989), querying one neighbor takes $\mathcal{O}\!\left(d\log n\right)$ time on average. In the worst case, query time is $\mathcal{O}\!\left(dn\right)$. One builds a ball-tree in $\mathcal{O}\!\left(dn\log n\right)$ time.
We construct a bpa $\operatorname{m}_{i}$ by storing its focal sets within a hash map and combine all $\operatorname{m}_{i}$-s as defined in (4). There are at most $\min\!\left(2^{k},2^{l}\right)$ focal sets of $\operatorname{\tilde{m}}$: Each $\operatorname{m}_{i}$ has at most two focal sets producing $2^{k}$ combinations and there are at most $2^{l}$ possible subsets of $\operatorname{\mathcal{Y}}$. We take the minimum as both are upper bounds. Looking up a focal set in the hash-map requires $\mathcal{O}\!\left(l\right)$ time. Extracting a prediction and the reject option requires $\mathcal{O}\!\left(l^{2}\right)$ time.
Combining the above yields a worst-case complexity of $\mathcal{O}\!\left(dkn+l\left(\min\!\left(2^{k},2^{l}\right)+\max\!\left(k,l\right)\right)\right)$. Since $k$ and $l$ are constant, the nearest-neighbor search dominates. The average total runtime of the search is $\mathcal{O}\!\left(dk\log n\right)$.
@@ B.1 Expected Probability Mass
Figure 2: Comparison of simulating and calculating the expected belief / probability mass of the true label $\tilde{y}$ and of the most frequently co-occurring label $\tilde{y}_{\operatorname{c}}$ for $k\in\left[100\right]$, $l=3$, $p_{1}=0.4$, $p_{2}=0.35$, and $p_{3}=0.25$.
@@ Appendix D Additional Experiments
This section contains additional experiments and their hyperparameters: Section D.1 provides experiments justifying Assumption 4.1, Section D.2 lists the values of all relevant hyperparameters, and the parameter sensitivity of our approach is discussed in Section D.3.
@@
Given a fixed instance $\tilde{x}$, Assumption 4.1 describes the label distribution of the neighbors’ candidate sets. Thereby, we assume that the true label $\tilde{y}$ dominates the neighborhood. Label $\tilde{y}_{\operatorname{c}}$ co-occurs most often with the true label $\tilde{y}$. Otherwise, we assume uniform noise among the remaining noise labels.
Figure 3 demonstrates that those assumptions are satisfied for most of the four real-world datasets (see Section 5.3). In most cases, the true label $\tilde{y}$ occurs most often in the neighborhood of instance $\tilde{x}$. There are one or two other labels with which $\tilde{y}$ is commonly confused, which we model by $\tilde{y}_{\operatorname{c}}$. Apart from that, all remaining noise labels are distributed uniformly.
@@ D.2 Hyperparameters
  • Rc (Feng et al., 2020): We use the same base models as mentioned above for Proden.
@@ D.3 Parameter Sensitivity
When increasing $k$, our method’s behavior on different datasets can be grouped into two groups. On the bird-song, mir-flickr, and yahoo-news datasets, increasing $k$ past ten neighbors also increases the amount of irrelevant labeling information from those neighbors. Therefore, our approach produces less confident predictions. At the same time, the MCC score of confident predictions remains at roughly the same level. This is because irrelevant labeling information from neighbors increases at most at the same rate as $k$. In contrast, on the lost and msrc-v2 datasets, the MCC score of confident predictions drops sharply at a certain point while the number of confident predictions decreases similarly. This is because irrelevant labeling information increases more rapidly than $k$: The decrease of confidence in predictions is slower than the increase of irrelevant candidate labels.

### v4

@@ 4.4 Runtime Complexity
We decompose the overall runtime of our approach in Algorithm 1 into $k$-times querying one nearest neighbor and creating its bpa (Lines 2–9), the cost of Yager’s rule (Line 10), as well as extracting predictions in Lines 11-12. Using the ball-tree data structure (Omohundro, 1989), querying one neighbor takes $\mathcal{O}\!\left(d\log n\right)$ time on average. In the worst case, query time is $\mathcal{O}\!\left(dn\right)$. One builds a ball-tree in $\mathcal{O}\!\left(dn\log n\right)$ time. We construct a bpa $\operatorname{m}_{i}$ by storing its focal sets within a hash map and combine all $\operatorname{m}_{i}$-s as defined in (4). There are at most $\min\!\left(2^{k},2^{l}\right)$ focal sets of $\operatorname{\tilde{m}}$: Each $\operatorname{m}_{i}$ has at most two focal sets producing $2^{k}$ combinations and there are at most $2^{l}$ possible subsets of $\operatorname{\mathcal{Y}}$. We take the minimum as both are upper bounds. Looking up a focal set in the hash-map requires $\mathcal{O}\!\left(l\right)$ time as the key length is variable. Extracting a prediction and the reject option then requires $\mathcal{O}\!\left(l^{2}\right)$ time. Combining the above yields a worst-case complexity of $\mathcal{O}\!\left(dkn+l\left(\min\!\left(2^{k},2^{l}\right)+\max\!\left(k,l\right)\right)\right)$. Since $k$ and $l$ are constant, the nearest-neighbor search dominates. The average total runtime of the search is $\mathcal{O}\!\left(dk\log n\right)$.
@@ 5 Experiments
Section 5.1 lists the tested methods, Section 5.2 shows our experimental setup, and Section 5.3 collects our main findings. Additional results and a description of all hyperparameters can be found in Appendix D.
@@ 5.1 Algorithms for Comparison
While many PLL algorithms exist (see Section 2), we focus on state-of-the-art methods commonly used in the literature. We consider ten methods: Pl-Knn (Hüllermeier & Beringer, 2005), Pl-Svm (Nguyen & Caruana, 2008), Ipal (Zhang & Yu, 2015), Pl-Ecoc (Zhang et al., 2017), Proden (Lv et al., 2020), Cc (Feng et al., 2020), Valen (Xu et al., 2021), Pop (Xu et al., 2023), CroSel (Tian et al., 2024), and Dst-Pll (our proposed method). Note that we decided to not consider PiCO (Wang et al., 2022) as it is only applicable to image data. In our experiments, however, we consider various image and non-image datasets as discussed in Section 5.2.
We choose the parameters of all methods as recommended by their respective authors and use the same base models for all of the neural network approaches. Appendix D.2 discusses the choices of all hyperparameters in more detail. As base models, we pick the LeNet architecture (LeCun et al., 1998) for the MNIST datasets and a $d$-300-300-300-$l$ MLP (Werbos, 1974) for all other datasets. For Pl-Knn and our method, we use variational auto-encoders (Kingma & Welling, 2014) to reduce the feature space dimensionality of the MNIST datasets and compute the nearest neighbors on the hidden representations. As our competitors do not provide a reject option, we use a threshold on their model’s confidences, that is, the maximum probability outputs, to evaluate the trade-off between the fraction and accuracy of non-rejected predictions.
@@
Pl-Knn (2005) 81.2 ($\pm$ 13.9) 75.8 ($\pm$ 11.1) 92.2 ($\pm$ 04.1) 84.8 ($\pm$ 07.2) 53.4 ($\pm$ 10.9)
Pl-Svm (2008) 62.3 ($\pm$ 16.3) 43.8 ($\pm$ 16.4) 67.5 ($\pm$ 09.8) 47.8 ($\pm$ 08.2) 39.1 ($\pm$ 09.6)
Ipal (2015) 79.3 ($\pm$ 17.1) 75.3 ($\pm$ 18.3) 92.9 ($\pm$ 04.3) 88.5 ($\pm$ 06.6) 58.7 ($\pm$ 09.8)
Pl-Ecoc (2017) 63.7 ($\pm$ 13.1) 66.5 ($\pm$ 12.5) 64.3 ($\pm$ 14.3) 51.7 ($\pm$ 10.7) 46.2 ($\pm$ 10.2)
Proden (2020) 81.6 ($\pm$ 14.0) 78.1 ($\pm$ 13.2) 93.9 ($\pm$ 04.4) 88.2 ($\pm$ 06.0) 64.2 ($\pm$ 08.2)
Cc (2020) 81.3 ($\pm$ 14.2) 78.8 ($\pm$ 13.6) 93.9 ($\pm$ 04.5) 89.6 ($\pm$ 05.7) 49.2 ($\pm$ 29.7)
Valen (2021) 79.7 ($\pm$ 15.3) 75.6 ($\pm$ 12.7) 91.8 ($\pm$ 04.2) 83.4 ($\pm$ 08.4) 63.6 ($\pm$ 09.7)
Pop (2023) 81.5 ($\pm$ 14.0) 78.1 ($\pm$ 13.1) 93.9 ($\pm$ 04.5) 88.1 ($\pm$ 06.0) 63.6 ($\pm$ 08.4)
CroSel (2024) 79.9 ($\pm$ 17.2) 78.4 ($\pm$ 14.5) 94.2 ($\pm$ 04.5) 88.9 ($\pm$ 06.7) 46.3 ($\pm$ 28.9)
Dst-Pll (ours) 80.8 ($\pm$ 14.1) 75.4 ($\pm$ 10.7) 92.0 ($\pm$ 04.2) 84.5 ($\pm$ 07.2) 52.0 ($\pm$ 11.6)
@@
Table 3: Fraction of rejects and non-rejected test-set accuracy of all methods with standard deviations across all experimental settings and five repetitions with different random seeds. To obtain the results, we tune the thresholds $\theta$ such that each competitor has a number of rejects that is comparable to that of our method.
Algorithms Fraction of rejects ($\pm$ std.) Non-rejected test accuracy ($\pm$ std.)
Pl-Knn (2005) 50.19 % ($\pm$ 20.98 %) 91.23 % ($\pm$ 10.12 %)
Pl-Svm (2008) 50.19 % ($\pm$ 20.98 %) 74.40 % ($\pm$ 19.77 %)
Ipal (2015) 50.19 % ($\pm$ 20.98 %) 83.52 % ($\pm$ 16.08 %)
Pl-Ecoc (2017) 50.19 % ($\pm$ 20.98 %) 73.92 % ($\pm$ 17.11 %)
Proden (2020) 50.19 % ($\pm$ 20.98 %) 94.03 % ($\pm$ 08.23 %)
Cc (2020) 50.19 % ($\pm$ 20.98 %) 90.13 % ($\pm$ 17.89 %)
Valen (2021) 50.19 % ($\pm$ 20.98 %) 86.81 % ($\pm$ 13.04 %)
Pop (2023) 50.19 % ($\pm$ 20.98 %) 94.02 % ($\pm$ 08.20 %)
CroSel (2024) 50.19 % ($\pm$ 20.98 %) 89.71 % ($\pm$ 18.67 %)
Dst-Pll (ours) 50.15 % ($\pm$ 21.16 %) 95.49 % ($\pm$ 07.01 %)
5.2 Experimental Setup
[…]
5.3 Results
@@
Table 3 shows the non-rejected test accuracy of all methods across all experimental settings for a fixed fraction of rejects. To obtain the results, we tune the confidence thresholds $\theta\in[0,1]$ such that each competitor rejects a similar number of instances as our proposed approach. Our approach uses the threshold $\Delta>0$, for which we have proved several desirable guarantees in Theorem 4.2. Our method achieves superior test-set accuracy on the non-rejected predictions.
Figure 1 shows the reject trade-off for varying confidence (0 to 1) and $\Delta_{\operatorname{\tilde{m}}}$ (-1 to 1) thresholds on the ecoli and KMNIST datasets with instance-dependent noise as well as on the msrc-v2 real-world dataset. The x-axes show the fractions of predictions that are rejected. The y-axes show the accuracies of predictions that are not rejected. The plots show (fraction of rejects, non-rejected test-set accuracy)-pairs corresponding to different settings of the thresholds. Most methods have monotonic growth: The more predictions are rejected, the more accurate are non-rejected predictions. Also, it is desirable to be close to the top-left corner of the plots as one wants to achieve high accuracy while rejecting as few predictions as possible. Note that the point (1, 1) cannot be observed as the test-set accuracy is undefined if all predictions are rejected.
Given a desired rejection rate $\hat{\operatorname{r}}(\Gamma_{g})$, Figure 1 also allows for numerically finding the appropriate value of $\lambda$ that minimizes ^$\hat{\operatorname{\mathcal{R}}}_{\lambda}(\Gamma_{g})$. Varying $\lambda$ in (1), while having $\hat{\operatorname{err}}(\Gamma_{g})$ and $\hat{\operatorname{r}}(\Gamma_{g})$ fixed, yields a straight line in Figure 1 showing all possible trade-offs.
Our method provides the significantly best trade-offs compared to competitors (Table 2). These results empirically demonstrate our method’s superiority in reliably rejecting unsure predictions.
@@ 6 Conclusions
When misclassification is costly, reject options provide a principled way of alleviating the consequences of incorrect predictions. In this work, we have presented a novel nearest-neighbor-based partial-label learning algorithm that can reject predictions if uncertain. We have demonstrated the desirable properties of the proposed reject option both from a theoretical (Theorem 4.2) and practical perspective. Our wide range of experiments showed the effectiveness of our classification rule and reject option on supervised datasets with added artificial noise and partially labeled real-world datasets.
Acknowledgments
We thank the anonymous reviewers for their constructive comments. This work was supported by the German Research Foundation (DFG) Research Training Group GRK 2153: Energy Status Data — Informatics Methods for its Collection, Analysis and Exploitation and by the KiKIT (The Pilot Program for Core-Informatics at the KIT) of the Helmholtz Association.
@@ Appendix A Proofs
This section collects our proofs. Appendix A.1 contains the proof of Theorem 4.2. The proof of Lemma 4.5 is in Appendix A.2. We prove our main statement, Theorem 4.6, in Appendix A.3.
A.1 Proof of Theorem 4.2
[…]
A.2 Proof of Lemma 4.5
@@ B.1 Expected Probability Mass
Figure 2: Comparison of simulating and calculating the expected belief of the true label $\tilde{y}$ and of the most frequently co-occurring label $\tilde{y}_{\operatorname{c}}$ for $k\in\left[100\right]$, $l=3$, $p_{1}=0.4$, $p_{2}=0.35$, and $p_{3}=0.25$.
@@ Appendix D Additional Experiments
This section contains additional experiments and details their hyperparameters. Section D.1 provides experiments justifying Assumption 4.4, Section D.2 lists the values of all relevant hyperparameters, the parameter sensitivity of our approach is discussed in Section D.3, and Section D.4 contains reject trade-off plots for all experimental settings considered.
@@
Given a fixed instance $\tilde{x}$, Assumption 4.4 describes the label distribution of the neighbors’ candidate sets. Thereby, we assume that the true label $\tilde{y}$ dominates the neighborhood and that label $\tilde{y}_{\operatorname{c}}$ co-occurs most often with the true label $\tilde{y}$. Otherwise, we assume uniform noise among the remaining noise labels.
Figure 3 demonstrates that those assumptions are largely satisfied on the four real-world datasets (see Section 5.3). In most cases, the true label $\tilde{y}$ occurs most often in the neighborhood of instance $\tilde{x}$. There are one or two other labels with which $\tilde{y}$ is commonly confused, which we model by $\tilde{y}_{\operatorname{c}}$. Apart from that, all remaining noise labels are distributed uniformly.
@@ D.2 Hyperparameters
  • CroSel (Tian et al., 2024): We use the same base models as mentioned above for Proden. We use 10 warm-up epochs using Cc and $\lambda_{cr}=2$. We abstain from using the data augmentations discussed in the paper for a fair comparison.
@@ D.3 Parameter Sensitivity
When increasing $k$, our method’s behavior on different datasets can be assembled into two groups. On the bird-song, mir-flickr, and yahoo-news datasets, increasing $k$ past ten neighbors also increases the amount of irrelevant labeling information from those neighbors. Therefore, our approach produces less confident predictions. At the same time, the MCC score of confident predictions remains at roughly the same level. This is because irrelevant labeling information from neighbors increases at most at the same rate as $k$. In contrast, on the lost and msrc-v2 datasets, the MCC score of confident predictions drops sharply at a certain point while the number of confident predictions decreases similarly. This is because irrelevant labeling information increases more rapidly than $k$: The decrease of confidence in predictions is slower than the increase of irrelevant candidate labels.
D.4 Reject Trade-off Curves
Figure 5 (i) – (xxxiv) shows the reject trade-off for varying confidence (0 to 1) and $\Delta_{\operatorname{\tilde{m}}}$ (-1 to 1) thresholds and augments Figure 1 by considering all datasets and noise generation strategies. The x-axes show the fractions of predictions that are rejected. The y-axes show the accuracies of predictions that are not rejected. The plots show (fraction of rejects, non-rejected test-set accuracy)-pairs corresponding to different settings of the thresholds. In most cases, our method provides a better trade-off between the number of rejected predictions and the accuracy of the non-rejected predictions. Table 2 summarizes all plots by showing the average empirical risks across all experimental settings and for different trade-off parameters $\lambda$. We recall that our method provides the significantly best trade-offs for $\lambda\in[0,0.2]$.
(i) ecoli (uniform noise)    (ii) ecoli (class-dep. noise)    (iii) ecoli (inst.-dep. noise)
(iv) m.-feats (uniform noise)    (v) m.-feats (class-dep. noise)    (vi) m.-feats (inst.-dep. noise)
(vii) pen-digits (uniform noise)    (viii) pen-digits (class-dep. noise)    (ix) pen-digits (inst.-dep. noise)
(x) semeion (uniform noise)    (xi) semeion (class-dep. noise)    (xii) semeion (inst.-dep. noise)
(xiii) solar-flare (uniform noise)    (xiv) solar-flare (class-dep. noise)    (xv) solar-flare (inst.-dep. noise)
(xvi) stat.-landsat (uniform noise)    (xvii) stat.-landsat (class-dep. noise)    (xviii) stat.-landsat (inst.-dep. noise)
(xix) theorem (uniform noise)    (xx) theorem (class-dep. noise)    (xxi) theorem (inst.-dep. noise)
(xxii) MNIST (uniform noise)    (xxiii) MNIST (class-dep. noise)    (xxiv) MNIST (inst.-dep. noise)
(xxv) FMNIST (uniform noise)    (xxvi) FMNIST (class-dep. noise)    (xxvii) FMNIST (inst.-dep. noise)
(xxviii) KMNIST (uniform noise)    (xxix) KMNIST (class-dep. noise)    (xxx) KMNIST (inst.-dep. noise)
(xxxi) bird-song                   (xxxii) mir-flickr
(xxxiii) yahoo-news                   (xxxiv) msrc-v2
Figure 5: Trade-off between the fraction of rejected predictions and the accuracy of non-rejected predictions.